# HierarchicalPrune: Position-Aware Compression for Large-Scale Diffusion Models

### v2

@@ 1 Introduction
Diffusion-based text-to-image (T2I) synthesis (2021ICLRPxTIG12RRHS_SDE; stability_ai_sdxl_turbo; 2022arXiv221002747L_FlowMatching; 2022arXiv220600364K_karras; 2022arXiv221209748P_dit; 2024arXiv240212376L_fit) has emerged as a powerful tool for a wide variety of applications, such as the generation of educational content, creative artwork, and UX/UI design prototyping, intensifying demand for deployable billion-parameter diffusion models (DMs). While recent advances, such as Stable Diffusion 3.5 (SD3.5) (esser2024mmdit) and FLUX (blackforestlabs_flux1), significantly outperform previous generations (SDXL (2023arXiv230701952P_sdxl), SD1.5 (2022CVPRLDM), and DALLE-2) in image quality and text alignment, they also come with excessive model sizes (8-11B parameters) and compute demands, limiting the accessibility of such advanced models.
At the same time, despite quantitative metrics, e.g., GenEval (ghosh2023geneval) and DPG-Bench (hu2024ellaequipdiffusionmodels), suggesting smaller models with approx. 2B parameters like SANA-Sprint (chen2025sanasprintonestepdiffusioncontinuoustime) perform better than their large-scale counterparts, user-based evaluations on the Artificial Analysis Leaderboard [^1] https://artificialanalysis.ai/image/leaderboard/text-to-image reveal a significant gap in the perceived quality between compact models and their larger counterparts, which quantitative metrics fail to capture. As such, there is a longstanding need for deploying large models even in resource-constrained settings in order to provide users with high-quality T2I capabilities.
@@
Nonetheless, this endeavour constitutes a very challenging task due to the high memory and compute intensity (li2025svdquant) of large DMs, thereby often requiring cloud solutions equipped with high-end GPUs and a minimum of 80GB VRAM. Concurrently, most of the renowned recent releases of foundational DMs, including SD3.5 (esser2024mmdit), FLUX (blackforestlabs_flux1), and Seedream 3.0 (2025seedream3), are built upon multi-modal diffusion Transformer (MMDiT) backbones (esser2024mmdit) instead of U-Net, revealing new compression opportunities (Section 2.1). At the moment, existing efforts to improve efficiency face important limitations. Firstly, sampling step reduction (2023arXiv230600980L_snapfusion; stability_ai_sdxl_turbo; 2023arXiv231117042S_add) and efficient operator design (xie2025sana; FlashAttention) are tailored to improving the speed of DMs rather than reducing memory requirements, leaving the important task of DM deployment in memory-constrained devices unresolved. Secondly, existing depth-pruning methods (lee2024koala; kim2024bksdm; fang2024tinyfusion) show promising results in both memory and computation reduction (kim2024bksdm; fang2024tinyfusion), outperforming width pruning (fang2023structural; Castells_2024_CVPR_ldpruner), but face critical scalability challenges. While they achieve reasonable compression on U-Net-based small DMs (2.6B or less), they fail to compress large-scale, state-of-the-art (SOTA) DMs such as SD3.5 Large (8B) and FLUX (11B) without experiencing significant degradation at 20-30% memory reduction as demonstrated in Table 1. The full-block-removal methods employed therein cannot capture the fine-grained impact of the subcomponents within each block, and more importantly, the different roles of blocks in different positions across the network’s hierarchy, leading to excessive performance drop at high pruning ratios. Lastly, orthogonal to pruning, reduced-precision computation (he2023ptqd; li2025svdquant; 2024arXiv240203666W_quest) has been proven effective, but it has not been combined with prior block-removal DM compression methods.
@@ 2.1 Motivation
Examining MMDiT-based DMs, we identify a two-fold hierarchy: i) inter-block and ii) intra-block hierarchy. Inter-block hierarchy refers to the organisation of blocks along the overall architecture and implies a functional hierarchy where blocks have different responsibilities based on their position. Intra-block hierarchy focuses on the fact that individual blocks are internally composed of subcomponents, which in the case of MMDiT consist of Norm, Context Norm, Attention, MLP, and Context MLP modules (esser2024mmdit), with each subcomponent having a varying impact on performance based on its type and position.
@@
To investigate this, we conducted a contribution analysis on SD3.5 Large Turbo over the HPSv2 dataset (Wu_2023_HPS), by removing both individual MMDiT blocks (Fig. 4a) and subcomponents (Fig. 4b4c and Fig. 8b8e in Appendix B), and comparing the performance before and after. We observe that each of the subcomponents, as well as whole-block removal, demonstrates different patterns of impact at different locations. This was further highlighted when analysing the joint removal of multiple subcomponent types, where different subcomponent combinations exhibited significantly different effects on the final performance (see Appendix B).
@@
Despite this variability, existing full-block removal methods (lee2024koala; kim2024bksdm) treat MMDiT blocks homogeneously, compromising performance when targeting high compression ratios. Our observations reveal a critical shortcoming: by ignoring the inter-block differences, these methods inadvertently prune blocks that are disproportionately important to visual quality while retaining less impactful ones. Moreover, by coarsely removing whole blocks, existing methods not only discard redundant layers but also eliminate subcomponents that might be essential for capturing fine-grain features. As shown in Section 3.2, this lack of differentiation leads to a steep drop in model performance under aggressive pruning rates. These insights motivate us to design hierarchy-informed techniques for the effective compression of large MMDiT-based DMs.
@@
Figure 4: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block (a), following prior depth pruning approaches (lee2024koala; kim2024bksdm; fang2024tinyfusion), or an intra-block subcomponent (b, c and see Fig. 8 in Appendix B for full set of analysis). We report the performance drop compared to the original model. The discrepancy in performance drop patterns reveals the different patterns of importance of each subcomponent.
@@ 3.1 Experimental Setup
Architectures, Datasets, and Implementation: We target SD3.5 Large Turbo (8B) and FLUX.1-Schnell (12B), two SOTA models designed to perform diffusion tasks within a small number of steps (e.g., 4). Following prior work (lee2024koala), we use the YE-POP dataset (yepop) consisting of 500K images, derived from LAION-5B (2022NeurIPSLAION5B). We implemented the HierarchicalPrune pipeline in PyTorch, building upon SD3.5 Large Turbo and FLUX.1-Schnell model checkpoints and pipelines from Diffusers (von-platen-etal-2022-diffusers). For 4-bit weight quantisation, we adopt bitsandbytes (dettmers2022llmint8; dettmers2023qlora). Note that other quantisation methods, such as SVDQuant (li2025svdquant), can also be combined with HierarchicalPrune, and we leave such exploration as future work. We aim to release our codebase upon acceptance of the paper.
@@
Baselines: We compare HierarchicalPrune with two prior works related to depth pruning and distillation: i) KOALA (lee2024koala), and ii) BK-SDM (kim2024bksdm). BK-SDM proposed block pruning of U-Net-based models using the CLIP score (hessel-etal-2021-clipscore), followed by distilling the pruned model using knowledge from the original model. KOALA follows a similar method, replacing CLIP-score-based importance ranking with scores from each block’s input-output cosine similarity. Also, we compare against SANA-Sprint-1.6B (chen2025sanasprintonestepdiffusioncontinuoustime), a SOTA compact model optimised for efficient on-device deployment.
Metrics: To quantitatively evaluate all methods, we employ the two most recent and representative image quality metrics, GenEval (ghosh2023geneval) (used in (xie2025sana; esser2024mmdit)) and HPSv2 (wu2023human) (used in KOALA (lee2024koala)). Furthermore, we conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics. Similar to (sheynin2023emueditpreciseimage) and (dai2023emuenhancingimagegeneration), we evaluate two aspects: (i) Text Alignment: How well a generated image follows the description of the text prompt, (ii) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors. On the system resource side, we report the measured peak memory usage and latency of running the target models on diverse GPUs, including A6000 (Table 2), GTX 3090 and A100 GPUs (see Appendix Table 5).
@@ 3.2 Main Results
Linear DiT SANA-Sprint 3.14 GB (100%) 0.77 29.61 -
Original 15.8 GB (100%) 0.71 30.29 -
KOALA 12.6 GB (79.4%) 0.37 19.99 41.2%
KOALA (+Quant) 3.56 GB (22.5%) 0.33 18.44 46.4%
SD3.5 BK-SDM 12.6 GB (79.4%) 0.38 21.21 38.2%
Large Turbo BK-SDM (+Quant) 3.56 GB (22.5%) 0.34 19.83 43.3%
Ours (HPP+PWP+Q) 3.56 GB (22.5%) 0.69 28.15 4.8%
Ours (All) 3.24 GB (20.5%) 0.62 26.29 13.3%
Original 22.6 GB (100%) 0.66 29.71 -
FLUX.1 KOALA 15.9 GB (70.5%) 0.38 25.24 28.7%
Schnell BK-SDM 15.9 GB (70.5%) 0.45 27.38 19.8%
Ours (All) 4.44 GB (19.6%) 0.64 28.69 3.2%
@@
Training Cost Comparison: Beyond quality preservation, HierarchicalPrune requires only 615–1,287 A100 GPU hours for architectural profiling and distillation (Appendix C.1), compared to 140k–200k A100 GPU hours required to pre-train small-scale DMs like SD1.4 and SD2.1 (2024ICLRgU58d5QeGv_cascade). This demonstrates that compressing existing large models is far more cost-efficient than training compact alternatives from scratch, while maintaining superior quality.
@@ 4 Related Work
T2I Diffusion Models. Since Stable Diffusion (SD) (2022CVPRLDM), there has been rapid community adoption and iterative updates in the field. SDv1.4 and v1.5 were released in 2022, enhancing efficiency and enabling specialised fine-tuning (e.g., DreamBooth). By mid-2023, Stable Diffusion XL (SDXL) significantly advanced resolution ($\times1024$), text comprehension, and image quality. By the end of 2024, the adoption of MMDiT (esser2024mmdit) as a backbone in SD3, SD3.5, and FLUX brought a significant boost in image generation quality and alignment with long text input. At the same time, the increasing parameter count, particularly in the backbone Transformer blocks, significantly improved image quality but with excessive resource demands. As SOTA models like SD3.5 scale up to as many as 8B parameters, efficient inference on resource-constrained devices becomes impractical. This underlines the significance of approaches such as our HierarchicalPrune framework, to enable DM deployment outside of high-end compute setups.
Compression Techniques. Traditional model compression methods, including distillation (hinton2015distilling), pruning (han2015deepcompression) (such as depth pruning via block removal (ghiasi2018dropblock; kim2024layermerge), width pruning (kwon2024tinytrain)), and quantisation (jacob2018quantization) have been applied to DMs (lee2024koala; kim2024bksdm; castells2024edgefusion; hu2024snapgen; fang2024tinyfusion; li2025svdquant; kim2024IDcompression; zhang2024laptopdiff), resulting in variable parameter efficiency gains. Specifically, (kim2024bksdm) and (lee2024koala) explore block removal in the U-Net backbone of SD1.5 and SDXL models, respectively. More recently, (li2025svdquant) explored the effectiveness of quantisation on DMs for memory reduction, hence our HierarchicalPrune employs adjunct post-training quantisation as part of its design. Orthogonal to compression, $\Delta$-DiT (chen2024deltaDit) identifies hierarchical patterns in DiT (MMDiT predecessor) for inference caching. Our analysis confirms that this generalises to modern MMDiT and further leverages these insights for the compression of SOTA multi-billion-scale DMs, under a unified framework.
@@ A.1 Datasets
We use the YE-POP dataset from HuggingFace for distillation training of the target models, similar to prior work by (lee2024koala). The YE-POP dataset is comprised of 500,000 high-quality images sampled from the LAION-5B dataset (2022NeurIPSLAION5B), incorporating a variety of visual modalities across multiple disciplines, including images of natural scenes, objects, people, and artistic images, providing extensive breadth for the training and evaluation of diffusion models (DMs). Our dataset preprocessing pipeline includes standard normalisation processes and low-resolution image and damage filtering, and ensures consistent image resolutions and aspect ratios to provide a fair data comparison for our model configurations and baseline methods.
@@ A.2 Architectures and Implementation
Implementation: The HierarchicalPrune pipeline is implemented in PyTorch. In particular, we built on top of checkpoints and inference pipelines for SD3.5 Large Turbo and FLUX.1-Schnell from the Diffusers library  (von-platen-etal-2022-diffusers). This allows us both consistency with pre- and post-processing steps, and also reproducible experiments. The distillation training process is conducted on a single node of 8 A100 GPUs with 80 GB of VRAM.
Quantization Strategy: To further reduce the memory footprint of the models, we employed 4-bit weight quantization from using bitsandbytes (dettmers2022llmint8; dettmers2023qlora).

### v3

@@ 1 Introduction
Diffusion-based text-to-image (T2I) synthesis (Song et al. 2021; Stability AI 2023; Lipman et al. 2023; Karras et al. 2022; Peebles and Xie 2022; Lu et al. 2024) has emerged as a powerful tool for a wide variety of applications, such as the generation of educational content, creative artwork, and UX/UI design prototyping, intensifying demand for deployable billion-parameter diffusion models (DMs). While recent advances, such as Stable Diffusion 3.5 (SD3.5) (Esser et al. 2024) and FLUX (Black Forest Labs 2024), significantly outperform previous generations (SDXL (Podell et al. 2023), SD1.5 (Rombach et al. 2022), and DALLE-2) in image quality and text alignment, they also come with excessive model sizes (8-11B parameters) and compute demands, limiting the accessibility of such advanced models.
At the same time, despite quantitative metrics, e.g., GenEval (Ghosh et al. 2023) and DPG-Bench (Hu et al. 2024b), suggesting smaller models with approx. 2B parameters like SANA-Sprint (Chen et al. 2025a) perform better than their large-scale counterparts, user-based evaluations on the Artificial Analysis Leaderboard [^1] https://artificialanalysis.ai/image/leaderboard/text-to-image reveal a significant gap in the perceived quality between compact models and their larger counterparts, which quantitative metrics fail to capture. As such, there is a longstanding need for deploying large models even in resource-constrained settings in order to provide users with high-quality T2I capabilities.
@@
Nonetheless, this endeavour constitutes a very challenging task due to the high memory and compute intensity (Li et al. 2025) of large DMs, thereby often requiring cloud solutions equipped with high-end GPUs and a minimum of 80GB VRAM. Concurrently, most of the renowned recent releases of foundational DMs, including SD3.5 (Esser et al. 2024), FLUX (Black Forest Labs 2024), and Seedream 3.0 (ByteDance 2025), are built upon multi-modal diffusion Transformer (MMDiT) backbones (Esser et al. 2024) instead of U-Net, revealing new compression opportunities (Section 2.1). At the moment, existing efforts to improve efficiency face important limitations. Firstly, sampling step reduction (Li et al. 2023; Stability AI 2023; Sauer et al. 2023) and efficient operator design (Xie et al. 2025; Dao et al. 2022) are tailored to improving the speed of DMs rather than reducing memory requirements, leaving the important task of DM deployment in memory-constrained devices unresolved. Secondly, existing depth-pruning methods (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024) show promising results in both memory and computation reduction (Kim et al. 2024a; Fang et al. 2024), outperforming width pruning (Fang et al. 2023; Castells et al. 2024a), but face critical scalability challenges. While they achieve reasonable compression on U-Net-based small DMs (2.6B or less), they fail to compress large-scale, state-of-the-art (SOTA) DMs such as SD3.5 Large (8B) and FLUX (11B) without experiencing significant degradation at 20-30% memory reduction as demonstrated in Table 1. The full-block-removal methods employed therein cannot capture the fine-grained impact of the subcomponents within each block, and more importantly, the different roles of blocks in different positions across the network’s hierarchy, leading to excessive performance drop at high pruning ratios. Lastly, orthogonal to pruning, reduced-precision computation (He et al. 2023; Li et al. 2025; Wang et al. 2024) has been proven effective, but it has not been combined with prior block-removal DM compression methods.
@@ 2.1 Motivation
Examining MMDiT-based DMs, we identify a two-fold hierarchy: i) inter-block and ii) intra-block hierarchy. Inter-block hierarchy refers to the organisation of blocks along the overall architecture and implies a functional hierarchy where blocks have different responsibilities based on their position. Intra-block hierarchy focuses on the fact that individual blocks are internally composed of subcomponents, which in the case of MMDiT consist of Norm, Context Norm, Attention, MLP, and Context MLP modules (Esser et al. 2024), with each subcomponent having a varying impact on performance based on its type and position.
@@
To investigate this, we conducted a contribution analysis on SD3.5 Large Turbo over the HPSv2 dataset (Wu et al. 2023b), by removing both individual MMDiT blocks (Fig. 4a) and subcomponents (Fig. 4b4c and Fig. 8b8e in Appendix B), and comparing the performance before and after. We observe that each of the subcomponents, as well as whole-block removal, demonstrates different patterns of impact at different locations. This was further highlighted when analysing the joint removal of multiple subcomponent types, where different subcomponent combinations exhibited significantly different effects on the final performance (see Appendix B).
@@
Despite this variability, existing full-block removal methods (Lee et al. 2024; Kim et al. 2024a) treat MMDiT blocks homogeneously, compromising performance when targeting high compression ratios. Our observations reveal a critical shortcoming: by ignoring the inter-block differences, these methods inadvertently prune blocks that are disproportionately important to visual quality while retaining less impactful ones. Moreover, by coarsely removing whole blocks, existing methods not only discard redundant layers but also eliminate subcomponents that might be essential for capturing fine-grain features. As shown in Section 3.2, this lack of differentiation leads to a steep drop in model performance under aggressive pruning rates. These insights motivate us to design hierarchy-informed techniques for the effective compression of large MMDiT-based DMs.
@@
Figure 4: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block (a), following prior depth pruning approaches (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024), or an intra-block subcomponent (b, c and see Fig. 8 in Appendix B for full set of analysis). We report the performance drop compared to the original model. The discrepancy in performance drop patterns reveals the different patterns of importance of each subcomponent.
@@ 3.1 Experimental Setup
Architectures, Datasets, and Implementation: We target SD3.5 Large Turbo (8B) and FLUX.1-Schnell (12B), two SOTA models designed to perform diffusion tasks within a small number of steps (e.g., 4). Following prior work (Lee et al. 2024), we use the YE-POP dataset (HuggingFace 2024) consisting of 500K images. We implemented the HierarchicalPrune pipeline in PyTorch and Diffusers (von Platen et al. 2022), utilising SD3.5 Large Turbo and FLUX.1-Schnell model checkpoints. For 4-bit weight quantisation, we adopt bitsandbytes (Dettmers et al. 2022).
@@
Baselines: We compare HierarchicalPrune with two prior works related to depth pruning and distillation: i) KOALA (Lee et al. 2024), and ii) BK-SDM (Kim et al. 2024a). BK-SDM proposed block pruning of U-Net-based models using the CLIP score (Hessel et al. 2021), followed by distilling the pruned model using knowledge from the original model. KOALA follows a similar method, replacing CLIP-score-based importance ranking with scores from each block’s input-output cosine similarity. Also, we compare against SANA-Sprint-1.6B (Chen et al. 2025a), a SOTA compact model optimised for efficient on-device deployment.
Metrics: To quantitatively evaluate all methods, we employ the two most recent and representative image quality metrics, GenEval (Ghosh et al. 2023) (used in (Xie et al. 2025; Esser et al. 2024)) and HPSv2 (Wu et al. 2023a) (used in KOALA (Lee et al. 2024)). Furthermore, we conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics. Similar to (Sheynin et al. 2023) and (Dai et al. 2023), we evaluate two aspects: (i) Text Alignment: How well a generated image follows the description of the text prompt, (ii) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors. On the system resource side, we report the measured peak memory usage and latency of running the target models on diverse GPUs, including A6000 (Table 2), GTX 3090 and A100 GPUs (see Appendix Table 5).
@@ 3.2 Main Results
Model Method Memory GenEval $\Uparrow$ HPSv2 $\Uparrow$ Reduction $\Downarrow$
Linear DiT SANA-Sprint 3.14 (100%) 0.77 29.61 -
SD3.5 Large Turbo Original 15.8 (100%) 0.71 30.29 -
KOALA 12.6 (79.4%) 0.37 19.99 41.2%
KOALA (+Quant) 3.56 (22.5%) 0.33 18.44 46.4%
BK-SDM 12.6 (79.4%) 0.38 21.21 38.2%
BK-SDM (+Quant) 3.56 (22.5%) 0.34 19.83 43.3%
Ours (HPP+PWP+Q) 3.56 (22.5%) 0.69 28.15 4.8%
Ours (All) 3.24 (20.5%) 0.62 26.29 13.3%
FLUX.1 Schnell Original 22.6 (100%) 0.66 29.71 -
KOALA 15.9 (70.5%) 0.38 25.24 28.7%
BK-SDM 15.9 (70.5%) 0.45 27.38 19.8%
Ours (All) 4.44 (19.6%) 0.64 28.69 3.2%
@@
Training Cost Comparison: Beyond quality preservation, HierarchicalPrune requires only 615–1,287 A100 GPU hours for architectural profiling and distillation (Appendix C.1), compared to 140k–200k A100 GPU hours required to pre-train small-scale DMs like SD1.4 and SD2.1 (Pernias et al. 2024). This demonstrates that compressing existing large models is far more cost-efficient than training compact alternatives from scratch, while maintaining superior quality.
@@ 4 Related Work
T2I Diffusion Models. Since Stable Diffusion (SD) (Rombach et al. 2022), there has been rapid community adoption and iterative updates in the field. SDv1.4 and v1.5 were released in 2022, enhancing efficiency and enabling specialised fine-tuning (e.g., DreamBooth). By mid-2023, Stable Diffusion XL (SDXL) significantly advanced resolution ($\times1024$), text comprehension, and image quality. By the end of 2024, the adoption of MMDiT (Esser et al. 2024) as a backbone in SD3, SD3.5, and FLUX brought a significant boost in image generation quality and alignment with long text input. At the same time, the increasing parameter count, particularly in the backbone Transformer blocks, significantly improved image quality but with excessive resource demands. As SOTA models like SD3.5 scale up to as many as 8B parameters, efficient inference on resource-constrained devices becomes impractical. This underlines the significance of approaches such as our HierarchicalPrune framework, to enable DM deployment outside of high-end compute setups.
Compression Techniques. Traditional model compression methods, including distillation (Hinton et al. 2015), pruning (Han et al. 2016) (such as depth pruning via block removal (Ghiasi et al. 2018; Kim et al. 2024c), width pruning (Kwon et al. 2024)), and quantisation (Jacob et al. 2018) have been applied to DMs (Lee et al. 2024; Kim et al. 2024a; Castells et al. 2024b; Hu et al. 2024a; Fang et al. 2024; Li et al. 2025; Kim et al. 2024b; Zhang et al. 2024), resulting in variable parameter efficiency gains. Specifically, (Kim et al. 2024a) and (Lee et al. 2024) explore block removal in the U-Net backbone of SD1.5 and SDXL models, respectively. More recently, (Li et al. 2025) explored the effectiveness of quantisation on DMs for memory reduction, hence our HierarchicalPrune employs adjunct post-training quantisation as part of its design. Orthogonal to compression, $\Delta$-DiT (Chen et al. 2024) identifies hierarchical patterns in DiT (MMDiT predecessor) for inference caching. Our analysis confirms that this generalises to modern MMDiT and further leverages these insights for the compression of SOTA multi-billion-scale DMs, under a unified framework.
@@ A.1 Datasets
We use the YE-POP dataset from HuggingFace for distillation training of the target models, similar to prior work by (Lee et al. 2024). The YE-POP dataset is comprised of 500,000 high-quality images sampled from the LAION-5B dataset (Schuhmann et al. 2022), incorporating a variety of visual modalities across multiple disciplines, including images of natural scenes, objects, people, and artistic images, providing extensive breadth for the training and evaluation of diffusion models (DMs). Our dataset preprocessing pipeline includes standard normalisation processes and low-resolution image and damage filtering, and ensures consistent image resolutions and aspect ratios to provide a fair data comparison for our model configurations and baseline methods.
@@ A.2 Architectures and Implementation
Implementation: The HierarchicalPrune pipeline is implemented in PyTorch. In particular, we built on top of checkpoints and inference pipelines for SD3.5 Large Turbo and FLUX.1-Schnell from the Diffusers library  (von Platen et al. 2022). This allows us both consistency with pre- and post-processing steps, and also reproducible experiments. The distillation training process is conducted on a single node of 8 A100 GPUs with 80 GB of VRAM.
Quantization Strategy: To further reduce the memory footprint of the models, we employed 4-bit weight quantization from using bitsandbytes (Dettmers et al. 2022, 2023).